OpenStudy (anonymous):

what represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (calculusxy):

I suggest you to use the formula: y=mx+b

OpenStudy (anonymous):

not before using \[y=y_1=m(x-x_1)\]

OpenStudy (anonymous):

The equation we'll be using here is: \[y-y_1=m(x-x_1)\]Where m is the slope of the line, and x_1,y_1 are the coordinates of any point on the line. We have the slope and a point on the line, so you can plug in and solve.

OpenStudy (anonymous):

oops i mean \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

I have no idea what you guys r saying